OpenStudy (anonymous):

Use the Pythagorean theorem to find one leg if the hypotenuse is 13 and the other leg is 5. Answer 12 3 13.93 11.24

OpenStudy (anonymous):

13^2-5^2 =169-25 =144 get the sqrt of 144 =12

OpenStudy (anonymous):

this is a right angled triangle worth remember sides 5,12,13 - called a pythagorean triad. because 13^2 = 12^2 + 5^2 - whole numbers fit the function. the lowest is 3,4 and 5. answer 12

OpenStudy (anonymous):

thank you both